a.) Find and identify the traces of the quadric surfacex^2+y^2-z^2= 1 and use the traces to classify the graph of the surface.
b.) If we change the equation in part a.) to x^2-y^2+z^2= 1, how is the graph affected?
c.) What if we change the equation in part a.) to x^2+y^2+ 2y-z^2= 0? How is the graph affected in this case?
1
Expert's answer
2020-09-14T19:20:27-0400

a) "x^2+y^2-z^2=1"


"x=k=> k^2+y^2-z^2=1=>y^2-z^2=1-k^2"

The trace is a hyperbola when "k\not=\pm 1."

If "k=\pm 1, y^2-z^2=(y-z)(y+z)=0," so it is a union of two lines.



"y=k=> x^2+k^2-z^2=1=>x^2-z^2=1-k^2"

The trace is a hyperbola when "k\not=\pm 1."

If "k=\pm 1, x^2-z^2=(x-z)(x+z)=0," so it is a union of two lines.



"z=k=> x^2+y^2-k^2=1=>x^2+y^2=1+k^2"

The trace is a circle whose radius is "\sqrt{1+k^2}."

Therefore the surface is a stack of circles, whose traces of other directions are

hyperbolas. So it is a hyperboloid. The intersection with the plane "z=k" is

never empty. This implies the hyperboloid is connected.


b) The role of "y" and "z" are interchanged. So now the axis of given hyperboloid

is "y" -axis.


c)


"x^2+y^2+2y-z^2=0=>"

"x^2+(y^2+2y+1)-z^2=1=>"

"x^2+(y+1)^2-z^2=1"

Thus it is a translation of the hyperboloid "x^2+y^2-z^2=1" by "(0,-1,0)."